# A Controllable and Realistic Framework for Evaluating Microservice Scheduling in Cloud-Edge Continuum

### v2

@@ 8.1. Cloud-edge Testbed Setup
Setup: We implemented and validated the proposed iDynamics framework using Kubernetes (open-source container orchestration system, 2024), the de facto container orchestration platform. The experimental cloud-edge cluster comprised one master node and nine worker nodes without preset anti-colocation rules such as node taints or pod affinity. The master node features 32 CPU cores (AMD EPYC 7763, x86_64 architecture), 32 GiB RAM, and 16 Gbps network bandwidth. Each worker node contains 4 CPU cores from the same AMD series, 32 GiB RAM, and 16 Gbps network bandwidth. The software stack includes Kubernetes v1.27.4, Calico v3.26.1 (CNI plugin), Istio v1.20.3 (service mesh), and CRI-O v1.27.1 as the container runtime. All nodes operate on Ubuntu 22.04.2 LTS with Linux kernel 5.15.0.
@@
Workload Generation: We adopted Social Network from DeathStarBench(Gan et al., 2019) as the microservice application workload and wrk2 (wrk2: An HTTP benchmarking tool based on wrk, 2024) as the requests generator to validate iDynamics and the proposed sample scheduling policies. Social Network benchmark emulates a simplified social media platform similar to popular social networking services. It is structured to replicate the intricate interactions and communication patterns in such types of applications. The benchmark comprises 27 different microservices that collectively offer request types such as composing a post, writing a user timeline, and writing a home timeline, which would trigger different call-graph topologies. wrk2 is capable of generating different types and proportions of workload requests for performance testing and measuring how well the cloud applications can respond to varying traffic. By combining the advantages of Social Network application and wrk2 benchmarking tool, various scenarios of workload generation can be achieved, thus facilitating the generation of call-graph dynamics and evaluation of different scheduling policies.
8.2. Effectiveness and Validation of Networking Dynamics Manager
Effectiveness: Precisely injecting and measuring diverse communication delays between nodes is challenging. We validated our delay injection method (depicted in Figure 6(b)) by injecting varied delays from a source node to multiple destinations and measuring the actual communication delays. Results in Table 1 demonstrate high accuracy, typically within 1 ms of the target delay, minor variations attributed to inherent randomness in cloud networking stacks. Additionally, to demonstrate that other traffic is not influenced, we tested the communication delays from all worker nodes to other destinations. One destination is the master node in the same cluster, and the other is the Google host (google.com), showing average delays of 0.50 $ms$ and 14.20 $ms$, respectively.
Similarly, bandwidth constraints were effectively enforced using our bandwidth shaping approach (illustrated in Figure 6(c)), as shown in Table 2. The measured bandwidth closely aligns with target values, confirming the efficacy of the method.
@@ 8.2. Effectiveness and Validation of Networking Dynamics Manager
8.3. Case Studies for Policy Evaluations
Table 3. Performance comparison between Kubernetes default scheduling and Policy 1 (Call-graph–Aware) under varying workloads and call-graph dynamics. Metrics include average (Avg), 99th percentile (p99), and standard deviation (Stdev) of response time (ms). SLA violations (¿150 ms) are highlighted in red, and SLA-compliant improvements are highlighted in green.
Policy Running time 0–8 min (Call-graph 1) 8–16 min (Call-graph 2) 16–24 min (Call-graph 3) 24–32 min (Call-graph 4, mixed)
QPS (req/sec) 30 10 50 70 30 10 50 70 30 10 50 70 30 10 50 70
K8s Policy Avg 13.27 14.36 13.22 13.22 168.6 171.8 171.2 174.2 57.33 56.06 49.92 46.11 40.67 44.15 39.31 40.15
p99 20.24 38.91 18.08 18.64 299.1 243.8 315.6 326.9 197.8 203.1 197.3 187.5 207.7 259.7 203.5 203.52
Stdev 2.46 4.61 2.51 1.88 24.87 20.17 33.17 39.46 48.3 45.73 38.64 31.42 51.43 60.34 49.63 50.49
Policy 1 Avg 13.89 14.84 13.51 13.33 168.88 127.6 117.26 118.4 71.32 70.87 68.4 61.83 34.92 39.37 35.2 35.69
p99 23.85 42.72 17.38 20.00 300.8 594.4 210.7 218.75 309.76 314.4 310.0 305.4 202.2 251.3 183.4 194.3
Stdev 3.26 5.26 2.10 2.13 25.27 92.80 19.12 22.33 68.42 68.38 63.62 49.84 43.76 51.00 43.75 43.88
8.3.1. Policy 1: Call-graph Scheduling Policy Evaluation
Figure 9. Average response time comparision between K8s-scheduling and Policy1 (Call-graph-Aware scheduling) under sustained and varying workloads.
In this case study, we evaluate the effectiveness of the Call-graph–Aware scheduling policy, which strategically places microservices with intensive inter-service communication onto nodes with high interconnectivity, reducing cross-node traffic.
@@ 8.3.1. Policy 1: Call-graph Scheduling Policy Evaluation
Discussion: Figure 9 and Table 3 clearly demonstrate Policy 1’s capability to rapidly mitigate SLA violations compared to the default Kubernetes scheduler. Upon SLA violation at the 10-minute point, Policy 1 effectively reduced the average response time from approximately 170 ms to about 120 ms, maintaining SLA compliance during heavy workloads. In contrast, Kubernetes default scheduling continued violating the SLA until the workload intensity decreased.
8.3.2. Policy 4: Hybrid-dynamics–Aware Policy Evaluation
Figure 10. Average response time comparision between Policy1 (Call-graph-Aware scheduling and Policy 4 (Hybrid-dynamics-Aware) under sustained workloads and changing networking conditions.
The Hybrid-dynamics–Aware policy integrates considerations of both call-graph dynamics and fluctuating network conditions (cross-node latency). It formulates microservice scheduling as a Service-Node Mapping Problem, aiming to jointly minimize communication costs arising from inter-service traffic volumes and network latencies.
@@ 8.3.2. Policy 4: Hybrid-dynamics–Aware Policy Evaluation
Discussion: As depicted in Figure 10, both policies maintained response times below the SLA target (¡ 300 ms). However, Policy 4 (Hybrid-dynamics–Aware) generally outperformed Policy 1 throughout the 40-minute evaluation. This confirms the Hybrid-dynamics–Aware policy’s enhanced ability to adapt dynamically and efficiently to combined network and QPS fluctuations, validating its suitability for complex and dynamic cloud-edge environments.
9. Conclusions and Future Work
In this paper, we presented iDynamics, a novel evaluation framework designed to systematically assess microservice scheduling policies under various dynamics in cloud-edge computing environments, particularly addressing call-graph dynamics and cross-node networking variability. The modular design of iDynamics enables realistic modelling and emulation of these dynamic conditions, along with providing extensible interfaces for users to develop and evaluate customized scheduling policies effectively.
For future work, we plan to extend the current framework by incorporating larger-scale clusters and heterogeneous resource provisioning capabilities. Additionally, we aim to explore and evaluate the performance of edge-intelligence workloads within a dynamically evolving cloud-edge continuum.

### v3

@@ VIII-A Cloud-edge Testbed Setup
Setup: We implemented and evaluated the proposed iDynamics framework on a Kubernetes‐based[36] cloud–edge testbed. The cluster consists of one control-plane node and fourteen worker nodes. The control-plane node has 32 CPU cores (AMD EPYC 7763, x86_64), 32 GiB of RAM, and a 16 Gbps network interface. Each worker node has 4 CPU cores from the same processor family, 32 GiB of RAM, and a 16 Gbps link. The software stack comprises Kubernetes v1.27.4 (deployed on ten nodes) and v1.28.4 (on five nodes), Calico v3.26.1 as the CNI plugin, Istio v1.20.3 as the service mesh, and CRI-O v1.27.1 (ten nodes) / v1.28.11 (five nodes) as the container runtime. All nodes run Ubuntu 22.04.2 LTS with the 5.15.0 Linux kernel.
@@
Workload Generation: We adopted Social Network from DeathStarBench[32] as the microservice application workload and wrk2 [37] as the requests generator to validate iDynamics and the proposed sample scheduling policies. Social Network benchmark emulates a simplified social media platform similar to popular social networking services. It is structured to replicate the intricate interactions and communication patterns in such types of applications. The benchmark comprises 27 different microservices that collectively offer request types such as composing a post, writing a user timeline, and writing a home timeline, which would trigger different call-graph topologies. wrk2 is capable of generating different types and proportions of workload requests for performance testing and measuring how well the cloud applications can respond to varying traffic. By combining the advantages of Social Network application and wrk2 benchmarking tool, various scenarios of workload generation can be achieved, thus facilitating the generation of call-graph dynamics and evaluation of different scheduling policies.
VIII-B Effectiveness and Validation of Networking Dynamics Manager
(a) 5 Cluster Nodes
(b) 10 Cluster Nodes
(c) 15 Cluster Nodes
Figure 10: Under varying call-graph dependencies, average response-time comparison between default Kubernetes scheduling and Policy 1 (Call-graph–Aware) under sustained and varying workloads for (a) 5, (b) 10, and (c) 15 cluster nodes.
(a) 5 Cluster Nodes
(b) 10 Cluster Nodes
(c) 15 Cluster Nodes
Figure 11: Under varying cross-node delays, average response-time comparison between Policy 1 (Call-graph–Aware scheduling) and Policy 4 (Hybrid-dynamics–Aware scheduling) under sustained workloads and changing networking conditions for (a) 5, (b) 10, and (c) 15 cluster nodes.
Effectiveness: Precisely injecting and measuring diverse communication delays between nodes is challenging. We validated our delay injection method (depicted in Figure 7b) by injecting varied delays from a source node to multiple destinations and measuring the actual communication delays. Results in Table II demonstrate high accuracy, typically within 1 ms of the target delay, with minor variations attributed to inherent randomness in cloud networking stacks. Additionally, we tested the communication delays from all worker nodes to other destinations to demonstrate that other traffic is not influenced. One destination is the master node in the same cluster, and the other is the Google host (google.com), showing average delays of 0.50 $ms$ and 14.20 $ms$, respectively.
Similarly, bandwidth constraints were effectively enforced using our bandwidth shaping approach (illustrated in Figure 7c), as shown in Table III. The measured bandwidth closely aligns with target values, confirming the method’s efficacy.
@@ VIII-B Effectiveness and Validation of Networking Dynamics Manager
VIII-C Two Case Studies for Policy Evaluations
In this subsection, we give the evaluations on three different cluster node scales (i.e., 5 cluster nodes, 10 cluster nodes, and 15 cluster nodes) and further discussions for two scheduling policies (i.e., Policy 1 and Policy 4) in subsection VII-C.
VIII-C1 Call-graph-Aware Scheduling Policy Evaluation
In this case study, we design, implement, and evaluate the effectiveness of a Call-graph–Aware scheduling policy, which strategically places microservices with intensive inter-service communication onto nodes with high interconnectivity, reducing cross-node traffic.
@@ VIII-C1 Call-graph-Aware Scheduling Policy Evaluation
Discussion: Figure 10 and Table IV demonstrate Policy 1’s capability to mitigate SLA violations compared to the default Kubernetes scheduler rapidly. For example, when evaluated with 10 cluster nodes, upon SLA violation at the 10-minute point, Policy 1 effectively reduced the average response time from approximately 170 ms to about 120 ms, maintaining SLA compliance during heavy workloads. In contrast, Kubernetes’ default scheduling continued violating the SLA until the workload intensity decreased. Similar effects of mitigating SLA violations from Policy 1 can also be observed in the evaluation scenarios of 5 cluster nodes and 15 cluster nodes.
VIII-C2 Hybrid-dynamics–Aware Policy Evaluation
The Hybrid-dynamics-aware policy integrates considerations of both call-graph dynamics and fluctuating network conditions (cross-node latency). It formulates microservice scheduling as a Service-Node Mapping Problem, aiming to jointly minimize communication costs arising from inter-service traffic volumes and network latencies.
@@ VIII-C2 Hybrid-dynamics–Aware Policy Evaluation
Discussion: As depicted in Figure 11, both policies respond to various SLA violations and maintain average response time below the SLA target ($<$300 ms). Under different cluster scales and varying cross-node delays, both policies can respond to SLA violations by adaptively scheduling the affected microservice applications, thus ensuring SLA guarantees.
IX Conclusions and Future Work
In this paper, we proposed iDynamics, a novel evaluation framework designed to systematically assess microservice scheduling policies under various dynamics in cloud-edge computing environments, particularly addressing call-graph dynamics and cross-node networking variability. The modular design of iDynamics enables realistic modelling and emulation of these dynamic conditions and provides extensible interfaces for users to develop and evaluate customized scheduling policies effectively. Experimental validation demonstrates the accuracy and effectiveness of our framework in capturing and reproducing intricate dynamics, thereby enabling reliable evaluation outcomes and insightful performance comparisons among competing scheduling strategies.
Although iDynamics provides significant advances in modelling dynamic cloud-edge environments, several important future research directions remain open. Firstly, the current implementation predominantly focuses on performance metrics such as traffic throughput and response time, yet energy consumption has become a critical consideration due to growing concerns about sustainability and operational costs. Therefore, we plan to incorporate comprehensive energy usage models into iDynamics, enabling policy evaluation under joint performance-energy optimization criteria.
Secondly, given the increasing emphasis on secure cloud-edge operations, future versions of iDynamics will integrate security constraints and mechanisms for policy enforcement and isolation. This extension aims to help researchers and practitioners analyze trade-offs between performance and security under realistic attack scenarios, enhancing the practical applicability of scheduling solutions.
Additionally, cloud-edge environments commonly involve multi-tenancy, which introduces complex interactions and resource contention among tenants. Our future work will extend iDynamics to include multi-tenant policy contention models, allowing detailed analyses of fairness, isolation, and interference management. Incorporating these aspects will make iDynamics an even more robust tool for evaluating scheduling policies in scenarios closely resembling real-world deployments, ultimately facilitating the development of adaptive, resilient, and efficient microservice scheduling strategies in the dynamic cloud-edge continuum.